OpenStudy (anonymous):

if x>2, then x^2-x-6 / x^2-4=

OpenStudy (anonymous):

Hi, Factor it. x² - x - 6 ----------- = x² - 4 (x - 3)(x + 2) ----------------- = (x - 2)(x + 2) Cancel out the x + 2 factors. x - 3 ------ <==ANSWER x - 2 I hope that helps!! :-)

OpenStudy (anonymous):

Thank You =)

OpenStudy (anonymous):

np